OpenStudy (anonymous):

Imagine a balance beam with unequal arms. An earring placed in the left basket was balanced by 4.71 g of standard masses on the right. When placed in the right basket, the same earring required 14.66 g on the left to balance. (a) Which was the longer arm? (b) What is the mass of the earring?

OpenStudy (ranga):

Let the left arm be 1 unit long and the right arm be x units long. Let M be the mass of the earring. When earring is placed on the left: M * 1 = 4.71 * x ----- (1) When earring is placed on the right: M * x = 14.66 * 1 ----- (2) Divide second equation by the first: x / 1 = 14.66 / (4.71x) x^2 = 14.66/4.71 = 3.1125 x = 1.76. Left arm is 1 unit and right arm is 1.76 units. So the right arm is longer A) The right arm is longer. B) Put x = 1.76 in (1): M = 4.71 * 1.76 = 8.31 grams.